three cards are drawn from a standard 52 card deck with replacement. What is the probability that at most 2 of the cards are diamonds?
Answer by stanbon(75887)

three cards are drawn from a standard 52 card deck with replacement. What is the probability that at most 2 of the cards are diamonds?
--------------------------------
Binomial Problem with n=3 and p(diamond) = 1/4
----
P(0<= x <=2) = 1 - P(x =3) = 1 - (1/4)^3 = 1 - 64 = 63/64
------------------------------------------------------------
Cheers,
Stan H.
===============